### v1

@@ 7.2. Functors Between Categories
Functors do exist if we focus on restrictions of $\mathbf{WMeas}_{\text{uc}}$ and $\mathbf{UHCont}$. In particular, call $\mathbf{WMeas}_{\text{uc,uhc}}\subset\mathbf{WMeas}_{\text{uc}}$ the category whose objects are measurable Polish spaces with the Borel $\sigma$-algebra, and whose morphisms are uniformly compact-valued, weakly measurable, and upper hemicontinuous correspondences. Then, the following holds.
@@ 7.3. Monads
The monad in Theorem 21 is the classical Vietoris monad (Garner,, 2020), adapted to the category of compact Hausdorff spaces with compact-valued, upper hemicontinuous correspondences.
@@
  • Miss part: $\nu_{X}^{-1}\langle U_{1},\dots,U_{m}\rangle=\bigl{\{}\mathcal{C}:\textstyle\bigcup_{C\in\mathcal{C}}C\subseteq U_{1}\cup\dots\cup U_{m},\;C\cap U_{i}\neq\emptyset\text{ }foralli\bigr{\}},$ again Vietoris-open.
@@
The monad that we found on the category $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$ is again the classic Vietoris monad, adapted to the category of compact Polish spaces (with their Borel $\sigma$-algebra), with uniformly compact-valued, weakly measurable correspondences. It is immediate to see that Theorem 8 still holds for $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$, since we simply require the extra assumption that the objects are compact measurable Polish spaces.
@@
    For every compact Hausdorff space $X$, pick any compatible complete separable metric $d_{X}$; The resulting topological space is compact Polish. Set $F(X)\coloneqq(X,d_{X})$ equipped with its Borel $\sigma$–algebra.
@@ 8. Conclusion
This work opens up the study of Categorical Conformal Prediction, and hints at many unsolved problems. For example, it is an open question how to leverage the structure of the functional-analysis-flavored categories UHCont and $\textbf{WMeas}_{\text{uc}}$ to learn further aspects of the CP methodology.
In addition, it is left to inspect whether studying CP as a morphism of different categories can allow us to remove the need for assumptions altogether, in particular compactness of the state-space $Y$.
@@ Acknowledgments
We wish to express our gratitude to Sam Staton and Paolo Perrone for insightful discussions on the possibility of formulating Conformal Prediction in the Category Theory framework. We are also extremely grateful to Sam Staton for his invitation to the University of Oxford, where the idea in the present paper germinated. Finally, we thank Yusuf Sale and Eyke Hüllermeier for stimulating discussions on the Uncertainty Quantification capabilities of Conformal Prediction, Sangwoo Park for his idea that led to Proposition 12, and Alessandro Zito for his patience in helping us shape the abstract and introduction.
@@ Appendix A Background in Functional Analysis
We say that $\Phi$ is upper hemicontinuous (uhc) if it is upper hemicontinuous at every point $x\in X$.
As the name suggests, an alternative notion of continuity for correspondences, called lower hemicontinuity, can be defined (Aliprantis and Border,, 2006, Definition 17.2). In this work, we focus primarily on uhc for its historical importance, especially in fields such as economics and game theory (Börgers,, 1991).

### v2

@@ 7.2. Functors Between Categories
Functors do exist if we focus on restrictions of $\mathbf{WMeas}_{\text{uc}}$ and $\mathbf{UHCont}$ [^12] We defer the study of monoidal functors to future research. In particular, call $\mathbf{WMeas}_{\text{uc,uhc}}\subset\mathbf{WMeas}_{\text{uc}}$ the category whose objects are measurable Polish spaces with the Borel $\sigma$-algebra, and whose morphisms are uniformly compact-valued, weakly measurable, and upper hemicontinuous correspondences. Then, the following holds.
@@ 7.3. Monads
The monad in Theorem 21 is the classical Vietoris monad (Garner,, 2020), adapted to the category of compact Hausdorff spaces with compact-valued, upper hemicontinuous correspondences. The Eilenberg-Moore category $\mathbf{AlgV}$ of algebras for the Vietoris monad on compact Hausdorff spaces is equivalent to the category $\mathbf{ContLat}$ of continuous lattices with Scott-continuous lattice homomorphisms (Day,, 1975; Wyler,, 1981).
@@
  • Miss part: $\nu_{X}^{-1}\langle U_{1},\dots,U_{m}\rangle=\bigl{\{}\mathcal{C}:\textstyle\bigcup_{C\in\mathcal{C}}C\subseteq U_{1}\cup\dots\cup U_{m},\;C\cap U_{i}\neq\emptyset\text{ }\forall i\bigr{\}},$ again Vietoris-open.
@@
The monad that we found on the category $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$ is again the classic Vietoris monad, adapted to the category of compact Polish spaces (with their Borel $\sigma$-algebra), with uniformly compact-valued, weakly measurable correspondences. It is immediate to see that Theorem 8 still holds for $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$, since we simply require the extra assumption that the objects are compact measurable Polish spaces. In addition, the same consideration regarding category $\mathbf{AlgV}$ of algebras for the Vietoris monad on compact Hausdorff spaces that we presented after Theorem 21, holds here as well, since Polish spaces are always Hausdorff.
@@
    For every metrizable compact Hausdorff space $X$, pick any compatible complete separable metric $d_{X}$; The resulting topological space is compact Polish. Set $F(X)\coloneqq(X,d_{X})$ equipped with its Borel $\sigma$–algebra.
@@ 8. Conclusion
This work opens up the study of Categorical Conformal Prediction, and hints at many unsolved problems. For example, it is an open question how to leverage the structure of the functional-analysis-flavored categories UHCont and $\textbf{WMeas}_{\text{uc}}$ to learn further aspects of the CP methodology. In particular, because the morphisms of these categories are correspondences, we are interested in what the Kleisli categories (Kleisli,, 1962, 1965), (Perrone,, 2024, Section 5.1) arising from the subcategories of $\mathbf{UHCont}$ and $\mathbf{WMeas}_{\text{uc}}$ admitting a monad, can tell us about Conformal Prediction.
In addition, it is left to inspect whether studying CP as a morphism of different categories can allow us to remove the need for assumptions altogether, in particular compactness of the state space $Y$.
@@ Acknowledgments
We wish to express our gratitude to Sam Staton and Paolo Perrone for insightful discussions on the possibility of formulating Conformal Prediction in the Category Theory framework. We are also extremely grateful to Sam Staton for his invitation to the University of Oxford, where the idea in the present paper germinated. We are indebted to Nicola Gambino for his assistance in improving the proof of Theorem 17, and for suggesting that we investigate the category of algebras for the Vietoris monad.
Finally, we thank Yusuf Sale and Eyke Hüllermeier for stimulating discussions on the Uncertainty Quantification capabilities of Conformal Prediction, Sangwoo Park for his idea that led to Proposition 12, and Alessandro Zito for his patience in helping us shape the abstract and introduction.
@@ Appendix A Background in Functional Analysis
We say that $\Phi$ is upper hemicontinuous (u.h.c.) if it is upper hemicontinuous at every point $x\in X$.
As the name suggests, an alternative notion of continuity for correspondences, called lower hemicontinuity, can be defined (Aliprantis and Border,, 2006, Definition 17.2). In this work, we focus primarily on u.h.c. for its historical importance, especially in fields such as economics and game theory (Börgers,, 1991).